A painter can paint a kitchen in 15 hours. An apprentice can paint the same kitchen in 30 hours. If they worked together, how long would it take them to paint the kitchen?
Answer by Cromlix(4381)

Painter can paint a kitchen in 15 hours
His apprentice can paint the same kitchen in 30 hours.
Painter's rate = 1/15 job/hr
Apprentice's rate = 1/30 job/hr
Together rate 1/x job/hr
1/15 + 1/30 = 1/x
Common Denominator = 30x
2x + x = 30
3x = 30
x = 10
They could paint the kitchen in 10 hrs together.
Hope this helps.
;-)